OpenStudy (anonymous):

Which is NOT a ratio comparing 2 to 10? a. 1/5 b. 3/10 c. 2/10 d. 2:10

OpenStudy (shamim):

U know 2:10=2/10=1/5

OpenStudy (shamim):

Ratio means 2:10

OpenStudy (shamim):

Ratio nd equal r same

OpenStudy (shamim):

We can divide by 2 both numerator nd denominator